OpenStudy (anonymous):

a sprinter starts from rest, and accelerates at 1.87m/s over a distance of 14.0m what is her final velocity?

OpenStudy (anonymous):

7.24

OpenStudy (dumbcow):

\[a(t) = 1.87\] \[v(t) = 1.87t\] \[x(t) = \frac{1.87}{2}t^{2}\] final velocity is at time when x(t) = 14 \[\frac{1.87}{2}t_f^{2} = 14\] \[t_f = \sqrt{\frac{28}{1.87}}\] \[v_f = 1.87*\sqrt{\frac{28}{1.87}}\]

OpenStudy (dumbcow):

yeah its about 7.236

OpenStudy (anonymous):

a car slows down at -5.00 m/s2 until it comes to stop after traveling 15.0m how much time did it take to stop? HELP ME!!!!!! PLEASE (:

OpenStudy (anonymous):

a jet fighter accelarates at 17.7 m/s2 increasing velocity from 119m/s to 233m/s. how much does that take?

OpenStudy (dumbcow):

both of these are found same way by defining acceleration , velocity, and distance functions \[a(t) = -5\] \[v(t) = -5t + v_o\] \[x(t) = -2.5t^{2}+v_o t\] car is stopped when velocity is 0 \[0 = -5t+v_o\] \[v_o = 5t\] car is stopped after 15m (x = 15) \[15 = -2.5t^{2}+(5t)*t\] solve for t \[2.5t^{2} = 15\] \[t = \sqrt{6}\]

OpenStudy (dumbcow):

2nd one, set up velocity function \[v(t) = 17.7t + v_o\] \[233 = 17.7t + 119\] solve for t \[t = \frac{233-119}{17.7}\]